# Neutral theory of cooperators

### v1

@@ 1 Introduction
In contrast, approach (ii) assumes that all species in the system are equivalent, shifting the focus to the effects of demographic noise as the main determinant of emerging ecological patterns [21, 22, 23, 24, 25]. Despite its simplicity, neutral theory successfully accounts for many relevant statistical patterns, including abundance distributions, species-area relations, and diversity estimates in space and time [26, 16, 22, 27]. However, a neutral theory that explicitly involves cooperative interactions is lacking, and little is known about the impact of mutualistic exchanges under a neutral picture. For example, microbiomes exhibit marked quantitative patterns that diverge significantly from those predicted by classic neutral theory, such as the presence of a persistent subset of species [12, 28, 29] or bimodal abundance distributions [30].
@@ 2 Model setup
Figure 1: Neutral model of cooperators (species indicated by colors). The population evolves according to two rules: (a) during replication, individuals of different species (black and red balls) cooperate. The red individual replicates and replaces a randomly chosen individual (blue ball). (b) during migration with probability $\mu$, an individual from a new species (yellow ball) enters from an external pool and substitutes a randomly-chosen individual (blue ball).
In the following, we consider the stationary properties of this stochastic process. We are particularly interested in the species abundance distribution $P_{n}$ (the probability that a species has abundance $n$ in steady state) and the system’s overall diversity. We measure diversity using the Simpson index [31], defined for abundance vector $\mathbf{n}$ as [^1] The Simpson index $\lambda(\mathbf{n})$ is sometimes alternatively defined as ${\Huge\Sigma}_{i=1}^{S}\frac{n_{i}(n_{i}-1)}{N(N-1)}$, i.e., as the probability of drawing two individuals of the same species when sampling without replacement. The definition (4) corresponds to the same probability when sampling with replacement. The two definitions differ by a small term of order $1/(N-1)$, which is irrelevant for our analysis.
@@ 3 Steady-state distribution
[…]
@@ 3.1.1 Low migration
Note that this approximation is only defined for $\mu\leq 1/\sqrt{2\pi N}$, so that the argument of the logarithm is nonnegative.
@@
Interestingly, an expression similar to Eq. (12) has previously appeared in a different ecological model, which considers a population embedded in a fluctuating spatiotemporal fitness landscape [39, Appendix C].
@@ 3.2 Steady-state number of species
With high migration, $R^{*}$ can be estimated by evaluating the denominator of (20). We may ignore the Gaussian contribution in (10), since the most relevant contribution to (10) is the Logseries distribution. For large $N$, the normalization constant of the Logseries term is $-\ln\mu$, thus we compute
@@ 4.1 Maximum abundances
u_1(ℓ) ≈[ e^ℓ^2/2N - (ℓ- 1)λ^* ( 2ND(x_ℓ) + 12 ) - 2ND(x_1) + 12 ]^-1,
[…]
@@ 5 Discussion
In this paper, we proposed a neutral model for cooperative ecosystems. As in the classic neutral theory of biodiversity [34], the strength of our approach lies in its minimalist design. Our analysis was divided into two parts, focusing on steady-state and dynamical properties, respectively.
@@
Several studies have reported bimodal abundance distributions in gut microbiomes, see for example Refs. [30, Fig. 1],  [40, Fig. 1], and [41, Fig. 2]. Previous research has attributed bimodality to emergent niche partitioning [42], intrinsic bistability [43], or a combination of multiple processes [41]. However, our theory suggests a neutral mechanism that generates bimodality under very minimal assumptions, offering a novel explanation of this phenomenon.
@@
In the second part of our paper, we studied the stochastic dynamics of species as they enter and leave the system. In the low-migration regime, the distribution of maximum abundances reached again exhibits a bimodal shape, allowing us to classify species into two types. This classification captures the dynamical signature of the cooperator core, as discussed in the steady-state analysis in the first part of this paper. We showed that species that enter the core achieve much higher abundances and longer residence times than species that do not enter the core.
Experimental validation of our predictions may be plausible using engineered microbial strains. For instance, by creating a synthetic consortium in which each species lacks the ability to synthesize one essential molecule, but is able to acquire it through cross-feeding with any other species. Such syntrophic communities have been successfully engineered, ranging from two-species [44, 45] to many [46].
@@
Further extensions of the theory should consider the heterogeneity of interactions, complex network topologies, and spatial effects. Another interesting direction would be to build a connection to the theory of autocatalytic chemical reaction networks, as studied in research on the origin of life. In particular, our model can be interpreted as a stochastic “hypercycle” [47, 48], i.e., a network of cross-catalytic polymers (or other chemical species) that replicate cooperatively. In this context, our results suggest a possible mechanism for the spontaneous emergence of stable and long-lived autocatalytic cores in chemical systems.
@@ Acknowledgements.
JP and RS thank the hospitality of the Santa Fe Institute, where this work started. JP also thanks the hospitality of the Simon Levin Lab and its members, where some of this work was developed. RS acknowledges support from the AGAUR 2021 SGR 0075 grant and the Santa Fe Institute. This project was partly supported by Grant No. 62417 from the John Templeton Foundation. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. AK was partly supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101068029. Finally, SR thanks NSF grant DMR-1910736 for financial support.
@@
Neutral theory of cooperators
@@ 1.1 Setup and notation
Due to our hypothesis of neutrality, all species evolve according to the same rules. Without loss of generality, we take $i=1$ to be the representative species, which, by neutrality, has the same statistical properties as any other species. It will be useful to define
$\mathbf{n}^{\circ}:=(n_{2},n_{3},\ldots)$ (S1.1)
as the vector of individual counts for the non-representative species. In order to alleviate notation, we interchangeably use $n_{1}\leftrightarrow n$ whenever appropriate.
A central concept of our analysis is the Simpson index, defined as
$\lambda(\mathbf{n})=\frac{\sum_{i}n_{i}^{2}}{\left(\sum_{i}n_{i}\right)^{2}}=\sum_{i}\left(\frac{n_{i}}{N}\right)^{2}.$ (S1.2)
where we used the fact that $\sum_{i}n_{i}=N$. As shown below, we also use the Simpson index for the non-representative species,
$\lambda(\mathbf{n}^{\circ})=\frac{\sum_{i>1}n_{i}^{2}}{\left(\sum_{i>1}n_{i}\right)^{2}}=\sum_{i>1}\left(\frac{n_{i}}{N-n_{1}}\right)^{2}$ (S1.3)
Finally, we denote the steady-state expectation of the Simpson index of all species as ${\lambda^{*}}$, and the Simpson index of the nonrepresentative species as $\lambda^{*}_{\circ}$.
@@ 1.2 Birth-death process describing the representative species
Box 1.1: A Box I.1: Birth probability. In a single timestep, the probability that the representative species undergoes the transition $n\to n+1$ is $(1-\mu)\left(\frac{n}{N}\right)\left(\frac{N-n}{N}\right)\left(\frac{N-n}{N}\right).$ (S1.4) Here, $1-\mu$ is the probability that the timestep involves a replication (rather than migration), $n/N$ is the probability that an individual from the representative species is chosen to replicate, $(N-n)/N$ is the probability that this individual is paired with an individual from a different species, and the final $(N-n)/N$ is the probability that the offspring replaces an individual of a different species. Since every migration event introduces a new species into the system, there is no contribution to the birth probability of the representative species due to migration.
Following from rules 1) and 2) in the main text, we derive the birth and death transition probabilities (see Boxes 1.2 and 1.2 for a full derivation) as
@@
$d_{n|\mathbf{n}^{\circ}}$ $=(1-\mu)\left(\frac{n}{N}\right)\left(1-\frac{n}{N}\right)\left[1-\left(1-\frac{n}{N}\right)\lambda({\mathbf{n}^{\circ}})\right]+\mu\left(\frac{n}{N}\right)$ (S1.6)
The birth-death process described by (S1.5)-(S1.6) is not closed. In particular, the death probability $d_{n|\mathbf{n}^{\circ}}$ depends on the Simpson index of the non-representative species, $\lambda(\mathbf{n}^{\circ})$. Assuming this Simpson index is tightly concentrated on its steady-state expectation value, $\lambda(\mathbf{n}^{\circ})\approx\lambda^{*}_{\circ}:=\langle\lambda_{\circ}(\mathbf{n})\rangle$, we may approximate steady-state fluctuations by a closed birth-death process:
$d_{n}$ $=(1-\mu)\left(\frac{n}{N}\right)\left(1-\frac{n}{N}\right)\left[1-\left(1-\frac{n}{N}\right)\lambda^{*}_{\circ}\right]+\mu\left(\frac{n}{N}\right).$ (S1.8)
Box 1.2: A Box I.2: Death transition probability. We first consider the probability that, in a given timestep, the representative species undergoes the transition $n\to n-1$ due to a cooperative interaction. Let $B$ and $C$ denote two species different from each other and from $A$ (the representative species). Then, we may have $(B+C)+A\rightarrow(B+C)+B\quad\text{or}\quad(B+A)+A\rightarrow(B+A)+B,$ where parentheses denote cooperative interaction. Adding up all possible contributions gives $\sum_{i\neq 1}\left(\frac{n_{i}}{N}\right)\left(\frac{N-n-n_{i}}{N}\right)\left(\frac{n}{N}\right)+\left(\frac{N-n}{N}\right)\left(\frac{n}{N}\right)\left(\frac{n}{N}\right)$ $=\left(\frac{N-n}{N}\right)\left(\frac{n}{N}\right)\sum_{i\neq 1}\left(\frac{n_{i}}{N}\right)-\left(\frac{n}{N}\right)\sum_{i\neq 1}\left(\frac{n_{i}}{N}\right)\left(\frac{n_{i}}{N}\right)+\left(\frac{n}{N}\right)\left(\frac{N-n}{N}\right)\left(\frac{n}{N}\right)$ $=\left(\frac{n}{N}\right)\left[\left(\frac{N-n}{N}\right)\left(\frac{N-n}{N}\right)-\sum_{i\neq 1}\left(\frac{n_{i}}{N}\right)\left(\frac{n_{i}}{N}\right)+\left(\frac{N-n}{N}\right)\left(\frac{n}{N}\right)\right]$ $=\left(\frac{n}{N}\right)\left(\frac{N-n}{N}\right)\left[1-\sum_{i\neq 1}\left(\frac{n_{i}}{N-n}\right)\left(\frac{n_{i}}{N-1}\right)\right]$ $=\left(\frac{n}{N}\right)\left(\frac{N-n}{N}\right)\left[1-\left(\frac{N-n}{N}\right)\lambda(\mathbf{n}^{\circ})\right],$ (S1.9) where in the last line we used definition (S1.3). In the death transition probability (S1.6), this sum is weighed by $1-\mu$, the probability that a replication event happens in a given timestep. In addition, the death probability includes the term $\mu(n/N)$, representing the probability $\mu$ that an individual of a new species enters the system times the probability $n/N$ that this new migrant replaces an individual of the representative species.

### v2

# Neutral theory of cooperative dynamics

@@ 1 Introduction
In contrast, approach (ii) assumes that all species in the system are equivalent, shifting the focus to the effects of demographic noise as the main determinant of emerging ecological patterns [21, 22, 23, 24, 25]. Despite its simplicity, neutral theory successfully accounts for many relevant statistical patterns, including abundance distributions, species-area relations, and diversity estimates in space and time [26, 16, 22, 27]. However, a neutral theory that explicitly involves cooperative interactions is lacking, and little is known about the impact of mutualistic exchanges under a neutral picture. For example, microbiomes exhibit marked quantitative patterns that diverge significantly from those predicted by classic neutral theory, such as the presence of a persistent subset of species [12, 28, 29].
@@ 2 Model setup
Figure 1: Neutral model of cooperators The population evolves according to two rules: (a) during replication, individuals of different species (black and red balls) cooperate. The red individual replicates and replaces a randomly chosen individual (blue ball). (b) during migration with probability $\mu$, an individual from a new species (yellow ball) enters from an external pool and substitutes a randomly-chosen individual (blue ball).
In the following, we consider the stationary properties of this stochastic process. We are particularly interested in the species abundance distribution $P_{n}$ (the probability that a species has abundance $n$ in steady state) and the system’s overall diversity. We measure diversity using the Simpson index [30], defined for abundance vector $\mathbf{n}$ as [^1] The Simpson index $\lambda(\mathbf{n})$ is sometimes alternatively defined as ${\Huge\Sigma}_{i=1}^{S}\frac{n_{i}(n_{i}-1)}{N(N-1)}$, i.e., as the probability of drawing two individuals of the same species when sampling without replacement. The definition (4) corresponds to the same probability when sampling with replacement. The two definitions differ by a small term of order 1/(N1)1/(N-1), which is irrelevant for our analysis.
@@ 3 Steady-state distribution
where $\langle\cdot\rangle$ indicates steady-state expectation over the entire population. Plugging (9) into (6) and using (8) we find that, under reasonable assumptions, the steady-state abundance distribution is approximated as (see SM2 for details):
@@ 3.1.1 Low migration
Note that this approximation is only defined for $\mu\leq 1/\sqrt{2\pi N}$, so that the argument of the square root is nonnegative.
@@
Interestingly, an expression similar to Eq. (12) has previously appeared in a different ecological model, which considers a population embedded in a fluctuating spatiotemporal fitness landscape [38, Appendix C].
@@ 3.2 Steady-state number of species
With high migration, $R^{*}$ can be estimated by evaluating the denominator of (20). We may ignore the Gaussian contribution in (10), since the most relevant contribution to (10) is the Logseries distribution. For large $N$, the normalization constant of the Logseries term is $\ln\mu$, thus we compute
@@ 4.1 Maximum abundances
$u_{1}(\ell)\approx\left[e^{\ell^{2}/2N-(\ell-1){\lambda^{*}}}\left(\sqrt{2N}\mathscr{D}(x_{\ell})+\frac{1}{2}\right)-\sqrt{2N}\mathscr{D}(x_{1})+\frac{1}{2}\right]^{-1},$
where we introduce the rescaled abundances $x_{\ell}:=(\ell-N{\lambda^{*}})/{\sqrt{2N}}$ and the Dawson function $\mathscr{D}(z):=e^{-z^{2}}\intop_{0}^{z}e^{t^{2}}dt$.
@@ 5 Discussion
In this paper, we introduced a neutral model for cooperative ecosystems, deliberately choosing the simplest set of rules consistent with cooperative interactions. This minimalist design, inspired by the neutral theory of biodiversity [33], allows a full analytical treatment while capturing the essential features of cooperative dynamics. Our analysis addresses both steady-state behavior and the dynamical properties of the system, providing a basic baseline for future comparisons with more complex models. It also complements existing models of cooperation, such as mutualistic voter models [39], where agents adopt neighbor states with a bias toward cooperation, or mutualistic Lotka-Volterra systems [40], where species benefit each other’s growth.
@@
Several studies have reported bimodal abundance distributions in gut microbiomes; see, for example, Refs. [41, Fig. 1] and [42, Fig. 2]. Previous research has attributed bimodality to emergent niche partitioning [43], intrinsic bistability [44], or a combination of multiple processes [42]. However, our theory suggests a neutral mechanism that generates bimodality under very minimal assumptions, offering a novel explanation of this phenomenon.
@@
In the second part of our paper, we studied the stochastic dynamics of species as they enter and leave the system. In the low-migration regime, the distribution of maximum abundances again exhibits a bimodal shape, allowing us to classify species into two types. This classification captures the dynamical signature of the cooperator core, as discussed in the steady-state analysis in the first part of this paper. We showed that species that enter the core achieve much higher abundances and longer residence times than species that do not enter the core.
It is worth noting that other models involving higher-order interactions have been proposed and investigated in various contexts. For example, Hinrichsen [45] investigated the universality class of the pair contact process $2A\to 3A$, $2A\to 0$ with diffusion and suggested that this model does not belong to the directed percolation universality class. In Ref. [46], the authors studied a quadratic contact process and determined a bound on the critical value of the infection rate, while in Refs. [47] and [48], the authors introduced general higher-order social contagion models and showed that these higher-order interactions fundamentally change the nature of the phase transition to a contagion phase. Although these models are related to ours in their use of higher-order interactions, they do not capture key features of our system, such as the bimodal species abundance distribution and the emergence of a cooperator core. Moreover, other similarities can be found when comparing our work with a recent Generalized Lotka-Volterra model with annealed disorder, which shows how temporal fluctuations in interactions act as effective environmental noise, promoting diversity [49]. Though based on different mechanisms, both highlight how non-equilibrium stochastic dynamics can generate high diversity and bimodal species abundance patterns.
In future work, it may be possible to generalize the method developed in this paper to other models of neutral cooperators, including three-way generalizations of our pairwise cooperation rule as well as other models considered in the literature. For a general neutral model, one may describe the abundance fluctuations of a “representative species” using a nonlinear birth-death master equation [34], with transition rates encoding mean-field effects of the rest of the population. Steady-state properties can then be identified by solving with self-consistency, in a similar way as done in this study.
Another promising direction to explore is the use of spatially extended models, which can offer additional insights into the effects of local interactions. This approach has also been extensively developed within the framework of neutral theory [33, 50, 25], where it has enabled the study of the dynamics of metacommunities. In our cooperative context, spatial structure would introduce correlations between neighboring organisms, effectively reducing the reproduction rate of existing species. We anticipate that this rate would scale with the surface area that separates a given species from its neighbors. Consequently, less frequent species—having a higher surface-area-to-volume ratio—would gain a relative advantage over more common species, which have a lower ratio. As a result, we expect that the mean abundance of the dominant core species would decrease in spatially structured environments.
Experimental validation of our predictions could be achieved using engineered microbial consortia. For instance, a synthetic community could be constructed in which each strain lacks the ability to synthesize a single essential metabolite but can obtain it through cross-feeding with the others. Such syntrophic systems have been successfully engineered, ranging from two-species [51, 52] to multispecies communities [53]. This approach would allow us to test our theoretical predictions under relaxed connectivity constraints, extending beyond the fully cooperative scenario assumed here. Minimal genetic modifications could be introduced into a single model organism (such as E. coli) to generate multiple auxotrophic variants that remain metabolically and physiologically equivalent, thus approximating a neutral interaction regime. By labeling some of these variants with a distinct fluorescent reporter, one could perform time-resolved sampling and quantify strain-specific abundance distributions using flow cytometry.
@@
Further extensions of the theory should consider the heterogeneity of interactions, complex network topologies, and spatial effects. Another interesting direction would be to build a connection to the theory of autocatalytic chemical reaction networks, as studied in research on the origin of life. In particular, our approach is related to the “hypercycle” [54, 55, 56, 57], a proposed model of collective chemical replication. In this model, a directed network of chemical species ($X_{i}$) cross-catalyze the replication of other species ($X_{i}+X_{j}\to X_{i}+X_{i}+X_{j}$ for neighbors $i\to j$). This reaction scheme is equivalent to our replication rule, Fig. 1, therefore our model can be interpreted as a (stochastic) hypercycle on a fully-connected network. In this context, our analysis could shed light on the emergence of cross-catalytic cores in replicating chemical systems.
@@ Acknowledgements.
JP and RS thank the hospitality of the Santa Fe Institute, where this work started. JP also thanks the hospitality of the Simon Levin Lab where some of this work was developed. RS acknowledges support from the Institució Catalana de Recerca i Estudis Avançats (ICREA), the AGAUR 2021 SGR 0075 grant, a PID2023-152129NB-I00 grant funded by MICIU/AEI/10.13039/501100011033 and by the Santa Fe Institute. This project was partly supported by Grant No. 62417 from the John Templeton Foundation. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. AK was partly supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101068029. Finally, SR thanks NSF grant DMR-1910736 for financial support.
@@
Neutral theory of cooperative dynamics
@@ 1.1 Setup and notation
Due to our hypothesis of neutrality, all species evolve according to the same rules. Without loss of generality, we take $i=1$ to be the representative species, which, by neutrality, has the same statistical properties as any other species. In order to alleviate notation, we interchangeably use $n_{1}\leftrightarrow n$ whenever appropriate.
A central quantity in our analysis is the Simpson index, defined as
$\lambda(\mathbf{n})=\frac{\sum_{i}n_{i}^{2}}{\left(\sum_{i}n_{i}\right)^{2}}=\sum_{i}\left(\frac{n_{i}}{N}\right)^{2}\,$, (S1.1)
where we used $\sum_{i}n_{i}=N$. Below, we also refer to the Simpson index of the non-representative species,
$\lambda_{\circ}(\mathbf{n})=\frac{\sum_{i>1}n_{i}^{2}}{\left(\sum_{i>1}n_{i}\right)^{2}}=\sum_{i>1}\left(\frac{n_{i}}{N-n_{1}}\right)^{2}\,.$ (S1.2)
Finally, we denote the steady-state expectation of the Simpson index of all species as $\lambda^{*}$}, and of the non-representative species as $\lambda^{*}_{\circ}$.
@@ 1.2 Birth-death process describing the representative species
Box 1.1: A Box I.1: Birth probability. In a single timestep, the probability that the representative species undergoes the birth transition $n\to n+1$ is $(1-\mu)\left(\frac{n}{N}\right)\left(\frac{N-n}{N}\right)\left(\frac{N-n}{N}\right).$ (S1.3) Here, $1-\mu$ is the probability that the timestep involves a replication (rather than migration), $n/N$ is the probability that an individual from the representative species is chosen to replicate, (Nn)/N(N-n)/N is the probability that this individual is paired with an individual from a different species, and the final (Nn)/N(N-n)/N is the probability that the offspring replaces an individual of a different species. Since every migration event introduces a new species into the system, there is no contribution to the birth probability of the representative species due to migration.
Following from rules 1) and 2) in the main text, we derive the birth and death probabilities (see Boxes 1.2 and 1.2 for a full derivation) as
@@
$d_{n}$ $=(1-\mu)\left(\frac{n}{N}\right)\left(1-\frac{n}{N}\right)\left[1-\left(1-\frac{n}{N}\right)\lambda_{\circ}({\mathbf{n}})\right]+\mu\left(\frac{n}{N}\right)\,.$ (S1.5)
The birth-death process described by (S1.4)-(S1.5) is not closed. In particular, the death probability $d_{n}$ depends on all the non-representative species abundances via their respective Simpson index, $\lambda_{\circ}(\mathbf{n})$. Assuming this Simpson index is tightly concentrated on its steady-state expectation value, $\lambda_{\circ}(\mathbf{n})\approx\lambda^{*}_{\circ}:=\langle\lambda_{\circ}(\mathbf{n})\rangle$, we may approximate steady-state fluctuations by a closed birth-death process:
$b_{n}$ $=(1-\mu)\left(\frac{n}{N}\right)\left(1-\frac{n}{N}\right)^{2},$ (S1.6)
$d_{n}$ $\approx(1-\mu)\left(\frac{n}{N}\right)\left(1-\frac{n}{N}\right)\left[1-\left(1-\frac{n}{N}\right)\lambda^{*}_{\circ}\right]+\mu\left(\frac{n}{N}\right),$ (S1.7)
where we note that $d_{n}$ now depends only on the abundance of the representative species $n$ and the expected $\lambda^{*}_{\circ}$.
Box 1.2: A Box I.2: Death probability. We first consider the probability that, in a given timestep, the representative species undergoes the death transition $n\to n-1$ due to a cooperative interaction. Let $B$ and $C$ denote two species different from each other and from AA (the representative species). Then, we may have $(B+C)+A\rightarrow(B+C)+B\quad\text{or}\quad(B+A)+A\rightarrow(B+A)+B,$ where parentheses denote cooperative interaction. Adding up all possible contributions gives $\sum_{i\neq 1}\left(\frac{n_{i}}{N}\right)\left(\frac{N-n-n_{i}}{N}\right)\left(\frac{n}{N}\right)+\left(\frac{N-n}{N}\right)\left(\frac{n}{N}\right)\left(\frac{n}{N}\right)$ $=\left(\frac{N-n}{N}\right)\left(\frac{n}{N}\right)\sum_{i\neq 1}\left(\frac{n_{i}}{N}\right)-\left(\frac{n}{N}\right)\sum_{i\neq 1}\left(\frac{n_{i}}{N}\right)\left(\frac{n_{i}}{N}\right)+\left(\frac{n}{N}\right)\left(\frac{N-n}{N}\right)\left(\frac{n}{N}\right)$ $=\left(\frac{n}{N}\right)\left[\left(\frac{N-n}{N}\right)\left(\frac{N-n}{N}\right)-\sum_{i\neq 1}\left(\frac{n_{i}}{N}\right)\left(\frac{n_{i}}{N}\right)+\left(\frac{N-n}{N}\right)\left(\frac{n}{N}\right)\right]$ $=\left(\frac{n}{N}\right)\left(\frac{N-n}{N}\right)\left[1-\sum_{i\neq 1}\left(\frac{n_{i}}{N-n}\right)\left(\frac{n_{i}}{N}\right)\right]$ $=\left(\frac{n}{N}\right)\left(\frac{N-n}{N}\right)\left[1-\left(\frac{N-n}{N}\right)\lambda_{\circ}(\mathbf{n})\right],$ (S1.8) where in the last line we used definition (S1.2). In the death probability (S1.5), this sum is weighted by $1-\mu$, the probability that a replication event happens in a given timestep. In addition, the death probability includes the term $\mu(n/N)$, representing the probability $\mu$ that an individual of a new species enters the system times the probability $n/N$ that this new migrant replaces an individual of the representative species.